OpenStudy (anonymous):

The hypotenuse of a right triangle is 20 centimeters, and the shorter leg is 12 centimeters. Find the length of the other leg.

OpenStudy (anonymous):

use pythagoras

OpenStudy (anonymous):

sounds like a quadruple-size 3-4-5 triangle to me!

OpenStudy (anonymous):

use the formula Other leg = Root (hypotenuse^2 - shortleg^2)

OpenStudy (anonymous):

a^2+b^2=c^2 so rearrange to be 20^2-12^2=b^2 which equals 400-144=256 square root it = 16

OpenStudy (anonymous):

20^2 = x^2 + 12^2

OpenStudy (anonymous):

other leg = root(20^2 - 12^2) = root (400 - 144) = root(256) = 16 so the other leg is 16cm